OpenStudy (anonymous):

What is the remainder when (x^3 – 58) ÷ (x – 4) ?

Parth (parthkohli):

Remainder Theorem: when \(p(x)\) is divided by \(x - k\), then the remainder is \(p(k)\).

mathslover (mathslover):

There are two methods for this 1) Long-division method 2) remainder theorem

OpenStudy (anonymous):

i dont recall the remainder theorem, care to explain?

mathslover (mathslover):

let x - 4 = 0 x = 4 p(x) = x^3-58 put x = 4 p(4) = 0 = 4^3 - 58

Parth (parthkohli):

Thus, you have to calculate \(f(4)\) where \(f(x) = x^3 - 58\)

OpenStudy (anonymous):

\[x=\sqrt[3]{58}\]

mathslover (mathslover):

\[\large{p(4) = 4^3-58=Remainder*}\] \[\large{64-58=Remainder}\]

mathslover (mathslover):

sorry I meant in my earlier post p(4) = remainder \(\ne 0\) ... @kakrazz hence we will get remainder = ?

OpenStudy (anonymous):

OH, okay thank you! I have never learned the remainder theorem. Thank you guys

OpenStudy (anonymous):

remainder is 6

OpenStudy (anonymous):

so 6

mathslover (mathslover):

very good @kakrazz good work

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you! i have like.. 8 more questions to ask on openstudy hahaha

mathslover (mathslover):

No problem kakrazz you can also use "Long-division" method ... BUT remainder theorem is best and easiest

OpenStudy (anonymous):

it seemed really easy too

mathslover (mathslover):

correct... it's easy but it's use is very wide